OpenStudy (anonymous):

WILL FAN AND MEDAL!!!!!what is the equation of the line that passes through the points (1 2/3) and (-1 -1)

OpenStudy (anonymous):

slope = y2-y1/x2-x1

OpenStudy (anonymous):

lets make x2, y2 = (-1, -1) and x1, y1 = (1/2, 3) then m = -1 -(3)/(-1-(1/2)) m= (-4/- 1.5) m=2.67 (approx)

OpenStudy (anonymous):

y- y1 = m(x - x1) y- 3 = 2.67 (x - 1/2)

OpenStudy (anonymous):

thanks! @BPDlkeme234 but I dont see that as one of the answers

OpenStudy (anonymous):

A. 2003-10-01-00-00_files/i0140002.jpg B. 2003-10-01-00-00_files/i0140003.jpg C. 2003-10-01-00-00_files/i0140004.jpg D. 2003-10-01-00-00_files/i0140005.jpg

OpenStudy (anonymous):

you dont have links to those

OpenStudy (anonymous):

http://learn.education2020.com/player/

OpenStudy (anonymous):

@BPDlkeme234